### v2

# Gromov-Hausdorff limits of smocked spaces

Abstract
Smocked spaces are a class of metric spaces which were introduced to generalise pulled thread spaces. We investigate convergence of these spaces, showing that the smocked space obtained from the Hausdorff limit of a sequence of smocking sets is equivalent to the Gromov-Hausdorff limit of the corresponding smocked spaces. We prove that it is sufficient to have uniform bounds on the smocking constants in order to have a sequence of smocked spaces with a Gromov-Hausdorff converging subsequence. We also show that given a normed space, there exists a smocked space whose unique tangent cone at infinity is that space. This answers three open questions of Sormani et al. We finish with some remarks on possible applications to metric measure spaces.
1 Introduction
The concept of a pulled thread space has existed for some time in metric geometry [1]. Roughly speaking, the idea is to take Euclidean space $\mathbb{E}^{N}$, view the space as a piece of cloth, sew along an interval or some other compact connected set in that space, and then pull the thread tight to obtain a new metric space where the sewing interval is now viewed as a point (a formal definition will be given shortly). Building on this concept, Sormani et al. recently introduced the notion of a smocked metric space and studied the balls and geodesics for several such spaces [2]. The definition is lengthy and will be given in Section 2, but one may think of a smocked space as a generalisation of a pulled thread space to include a (possibly countable) collection of disjoint compact connected sets on $\mathbb{E}^{N}$, all of which are ’pulled tight’.
One can prove that these spaces are well-defined as metric spaces and meet all the relevant axioms, but there remains a number of basic questions about convergence of these metric spaces and their subsequences, which we will address in this article [2]. One of the immediate applications mentioned in [2] is to apply results on Sormani-Wenger intrinsic flat convergence of smocked spaces to an open question of Gromov and Sormani on tori of almost nonnegative scalar curvature [3, 4]. In this article, we will not consider intrinsic flat convergence and will instead prove some rigorous results only for Gromov-Hausdorff convergence. It would certainly be desirable to have some more results of this kind before embarking on further studies of smocked spaces and their tangent cones at infinity [5]. Although a variety of smocked spaces and their tangent cones were studied in [2], it was necessary to carefully consider the explicit formulae for the distance functions in each case. In this article, we wish to show that it is possible to do analysis and geometry in smocked spaces in a general sense without having to know the distance function a priori.
The article is structured as follows. In Section 2, we state basic definitions and results. In Section 3, we prove that if one defines a sequence of smocked spaces using a sequence of smocking sets which has Hausdorff convergence to a limiting smocking set, then the corresponding smocked spaces converge in the pointed Gromov-Hausdorff sense. We also investigate sufficient conditions for a sequence of smocked spaces to have a subsequence which Gromov-Hausdorff converges, finding that it is sufficient to have uniform bounds on the smocking constants. Furthermore, we show that given a finite-dimensional normed space, there exists a smocked space with a unique tangent cone at infinity which is that space. Section 4 is somewhat more conjectural. We consider the possibility of extending to smocked spaces equipped with measures and also smocked sets which are defined for Wasserstein spaces of measures.
2 Preliminaries
2.1 Gromov-Hausdorff convergence
In this subsection, we present basic definitions concerning limits with the Gromov-Hausdorff distance.
Definition 2.1: Given two metric spaces $X$ and $Y$, the Gromov-Hausdorff distance is defined to be
$d_{GH}(X,Y)=\text{inf}\bigg{\{}d_{H}^{Z}(f(X),f(Y))\bigg{\}}$
taken over all metric spaces $Z$ and isometric embeddings $f:X\rightarrow Z$ and $g:Y\rightarrow Z$. $d_{H}^{Z}$ is the Hausdorff distance between subsets of $Z$
$d_{H}^{Z}(A,B)=\text{inf}\{\epsilon>0:B\subset T_{\epsilon}(A)\>\text{and}\>A\subset T_{\epsilon}(B)\},$
where $T_{\epsilon}(A)=\{x\in Z:d_{Z}(x,A)<\epsilon\}$.
Definition 2.2: If one has a sequence of complete noncompact metric spaces $(X_{j},d_{j})$ and points $x_{j}\in X_{j}$, pointed Gromov-Hausdorff convergence is defined as follows:
$(X_{j},d_{j},x_{j})\xrightarrow{pGH}(X_{\infty},d_{\infty},x_{\infty})$
if and only if for every $R>0$ the closures of the balls of radius $R$ in $X_{j}$ converge in the Gromov-Hausdorff sense as metric spaces with the restriction of the distance function to closures of balls in $X_{\infty}$
$(\overline{B}_{R}(x_{j}),d_{j})\xrightarrow{GH}(\overline{B}_{R}(x_{\infty}),d_{\infty}).$
A sequence of metric spaces can be obtained by repeatedly rescaling a space. Although some spaces like the taxi space are isometric to their rescalings, this is generally not the case. If a sequence or subsequence obtained by rescaling converges in the Gromov-Hausdorff sense, the limit space obtained is known as the tangent cone at infinity.
Definition 2.3: A complete non-compact metric space $(X,d_{X})$ with infinite diameter has a tangent cone at infinity $(Y,d_{Y})$ if there is a sequence of rescalings $R_{j}\rightarrow\infty$ and points $x_{0}\in X$ and $y_{0}\in Y$ such that
$(X,d/r_{j},x_{0})\xrightarrow{pGH}(Y,d_{Y},y_{0}).$
2.2 Smocked spaces
This section contains definitions and main results regarding smocked spaces.
Definition 2.4 Start with a finite or countable collection of disjoint connected compact sets in $\mathbb{E}^{N}$ called smocking stitches
$I=\{I_{j}:j\in J\}.$
Also, assume the existence of a positive smocking separation factor $\delta$
$\delta=\text{min}\{|z-z^{\prime}|:z\in I_{j},z^{\prime}\in I_{j^{\prime}},j\neq j^{\prime}\in J\}>0.$
Using these one can define a smocked metric space $(X,d)$ where each stitch in the collection is viewed as a single point:
$X=\{x:x\in\mathbb{E}^{n}\>\backslash\>S\}\cup I,$
where the smocking set $S$ is
$S=\bigcup_{j\in J}I_{j}.$
We also have a smocking map $\pi:\mathbb{E}^{N}\rightarrow X$ defined by
$\pi(x)=\left\{\begin{array}[]{ll}x&\mbox{if }x\in\mathbb{E}^{n}\>\backslash\>S\\ I_{j}&\mbox{if }x\in I_{j},\>j\in J\\ \end{array}\right.$
for some index set $J$ and a smocked distance function $d:X\times X\rightarrow[0,\infty)$ which is defined for $x,y\notin\pi(S)$ and stitches $I_{m}$ and $I_{n}$
$d(x,y)=\text{min}\{d_{0}(x,y),d_{1}(x,y),...\},$
$d(x,I_{n})=\text{min}\{d_{0}(x,z),d_{1}(x,z),...:z\in I_{n}\},$
$d(I_{m},I_{n})=\text{min}\{d_{0}(z^{\prime},z),d_{1}(z^{\prime},z),...:z^{\prime}\in I_{m},z\in I_{n}\},$
where $d_{k}$ is the sum of lengths of segments which go between $k$ stitches:
$d_{k}(v,w)=\text{min}\{|v-z_{1}|+\sum_{i=1}^{k-1}|z^{\prime}_{i}-z_{i+1}|+|z^{\prime}_{k}-w|:z_{i},z^{\prime}_{i}\in I_{j_{i}},j_{1}\neq...\neq j_{k}\in J\},$
for any $v,w\in\mathbb{E}^{n}$. The smocking pseudometric $\overline{d}:\mathbb{E}^{N}\times\mathbb{E}^{N}\rightarrow[0,\infty)$ is defined as
$\overline{d}(v,w)=d(\pi(v),\pi(w))=\text{min}\{d_{k}(v^{\prime},w^{\prime}):\pi(v)=\pi(v^{\prime}),\pi(w)=\pi(w^{\prime}),k\in\mathbb{N}\}.$
The definition of a smocked space refers to a set of four constants, which are referred to in [2] as the smocking constants: these are the smocking depth $h$, the maximum smocking length $L_{\text{max}}$, the minimum smocking length $L_{\text{min}}$, and the smocking separation factor $\delta$, which was already defined in Definition ?.
Definition 2.5 The smocking depth $h$ is defined as
$h=\text{inf}\{r:\forall x\in X\>\exists j\in J\>\exists z\in I_{j}\>\text{suchthat}\>|x-z|<r\}.$
Definition 2.6 The smocking lengths $L_{\text{max}}$ and $L_{\text{min}}$ are defined using the diameters of the smocking stitches as follows:
$L_{\text{min}}=\text{inf}\{D(I_{j}):\>j\in J\},$
$L_{\text{max}}=\text{sup}\{D(I_{j}):\>j\in J\}.$
The main theorem proved in [2] for smocked spaces is the following result:
Proposition 2.7: Given an $N$-dimensional smocked space $(X,d)$ such that
$|\overline{d}(x,x^{\prime})-|F(x)-F(x^{\prime})||\leq K,\>\>\>\>\>\>\forall\>x,x^{\prime}\in\mathbb{E}^{N},$
where $F:\mathbb{E}^{N}\rightarrow[0,\infty)$ is a norm, then $(X,d)$ has a unique tangent cone at infinity $(\mathbb{R}^{N},d_{F})$, where
$d_{F}(x,x^{\prime})=||x-x^{\prime}||_{F}=F(x,x^{\prime}).$
2.3 Metric geometry
The proofs of the results in Section 3 will require the definition of a correspondence for two metric spaces $X$ and $Y$:
Definition 2.8: Given two metric spaces $X$ and $Y$, a correspondence $C$ is a set $C\subset X\times Y$ such that
$\forall x\in X\>\exists y\in Y\>\text{such that}\>(x,y)\in C\>\text{and}\>\forall y\in Y\>\exists x\in X\>\text{such that}\>(x,y)\in C.$
Some of the proofs will also require the definition of an $\epsilon$-net.
Definition 2.9 An $\epsilon$-net in a pointed metric space $(X,x)$ is a subset $L$ of $X$ containing $x$ whose $\epsilon$-neighbourhood covers $X$ and for which there is some $\epsilon^{\prime}>0$ with $d(l_{1},l_{2})\geq\epsilon^{\prime}$ for all $l_{1}\neq l_{2}$ in $L$.
To any metric space $X$, one may also canonically associate another metric space called the Wasserstein space [14]. These spaces will be mentioned in Section 4.
Definition 2.10: The Wasserstein space $\mathcal{P}_{p}(X):=(P(X),W_{p})$ is a metric space associated to $X$, where $P(X)$ is the space of Borel probability measures on $X$ and $W_{p}$ is the Wasserstein distance of order $p$ between two measures $\mu$ and $\nu$, which is defined to be
[…]
where $\pi$ can take values in the set of all transference plans between $\mu$ and $\nu$.
3 Gromov-Hausdorff convergence of smocked spaces
3.1 Relation between Hausdorff and Gromov-Hausdorff convergence
We start with a result which relates Hausdorff convergence of smocking sets to Gromov-Hausdorff convergence of the corresponding smocked spaces.
Proposition 3.1: If one defines a sequence of smocked spaces $X_{k}$ using smocking sets $S_{k}\subset\mathbb{E}^{N}$ such that the sets Hausdorff converge in the limit to a smocking set $S_{\infty}$, then the corresponding smocked spaces converge to $X_{\infty}$ in the pointed Gromov-Hausdorff sense.
Proof: Smocked spaces are complete and noncompact (see Theorem 6.1 of [2]), so we use pointed Gromov-Hausdorff convergence as defined in Section 2. We must show that for all $r>0$
$\lim_{k\rightarrow\infty}d_{GH}\big{(}\big{(}B_{r}(x_{0}^{(k)}),d_{X_{k}}\big{)},\big{(}B_{r}(x_{0}^{(\infty)}),d_{X_{\infty}}\big{)}\big{)}=0,$
where $X_{\infty}$ is the smocked space obtained from the Hausdorff limit $S_{\infty}$ and $x_{0}^{(k)}$ is a base point in $X_{k}$.
Consider the smocking maps $\pi_{k}:\mathbb{E}^{N}\rightarrow X_{k}$ and $\pi_{\infty}:\mathbb{E}^{N}\rightarrow X_{\infty}$ associated to $X_{k}$ and $X_{\infty}$, respectively. By shifting the smocking set $S_{k}$, we can assume that
$\pi_{k}(0)=x_{0}^{(k)}\>\>\>\>\>\>\text{for any}\>\pi_{k},$
with a similar statement holding for $S_{\infty}$. We then define a subset $U_{r}(x_{0}^{(k)})$ of $\mathbb{E}^{N}$ defined for any $r>0$:
$U_{r}(x_{0}^{(k)})=\pi_{k}^{-1}(B_{r}(x_{0}^{(k)}))\subset\mathbb{E}^{N}.$
As our correspondence for each $r>0$, we take
$C=\{(\pi_{k}(w),\pi_{\infty}(w)):w\in U_{r}(x_{0}^{(k)})\}\subset B_{r}(x_{0}^{(k)})\times B_{r}(x_{0}^{(\infty)}).$
This is a correspondence because both smocking maps are surjections. Next, note that
$|d_{X_{k}}(\pi_{k}(v),\pi_{k}(w))-d_{X_{\infty}}(\pi_{\infty}(v),\pi_{\infty}(w))|=|\overline{d}_{k}(v,w)-\overline{d}_{\infty}(v,w)|.$
The collection of sets $S_{\infty}$ is a Hausdorff limit, so it is the collection of sets of limits of all converging sequences $\{s_{k}\}$ in $\mathbb{E}^{N}$ such that $s_{k}\in S_{k}$ for all $k$. In particular, such sequences can only grow with increasing $k$ with at most exponential speed, so the distances measured by the respective smocking pseudometrics can only decay with at most exponential speed. This implies that we have for the difference of the distances
$|\overline{d}_{k}(v,w)-\overline{d}_{\infty}(v,w)|\leq Ae^{-f(k)},$
where $A$ is some sufficiently large constant and $f(k)$ is a monotonically increasing function of $k$. On the other hand, such sequences could instead decay with at most exponential speed, and the same conclusion holds after taking the magnitude of the difference. By Theorem 7.3 of [2] for almost distance preserving correspondences, this implies that
$d_{GH}\big{(}\big{(}B_{r}(x_{0}^{(k)}),d_{X_{k}}\big{)},\big{(}B_{r}(x_{0}^{(\infty)}),d_{X_{\infty}}\big{)}\big{)}\leq 2Ae^{-f(k)}.$
Taking the limit, we have
$\lim_{k\rightarrow\infty}d_{GH}\big{(}\big{(}B_{r}(x_{0}^{(k)}),d_{X_{k}}\big{)},\big{(}B_{r}(x_{0}^{(\infty)}),d_{X_{\infty}}\big{)}\big{)}=0.$
[…]
Some explicit examples may help to illustrate this proposition:
Example 3.2: Start with $\mathbb{E}^{N}$ and take the sequence of smocking sets $S_{k}$ to be the sequence of closed balls about the origin with radius $1/k$. In the limit, $S_{\infty}$ is the point at the origin, so $X_{\infty}=\mathbb{E}^{N}$, since we have removed a point and replaced it with a point. Now for the sequence of smocked spaces, consider $X_{k}$:
$X_{k}=\{x\in\mathbb{E}^{N}\backslash\overline{B}_{1/k}(0)\}\cup\{\overline{B}_{1/k}(0)\}.$
In this space $\overline{B}_{1/k}(0)$ is identified with a point, so the distance between two points $x$ and $y$ is the minimum of the direct distance between $x$ and $y$ and all the possible distances between $x$ and $z\in\overline{B}_{1/k}(0)$ plus the distances between $z^{\prime}\in\overline{B}_{1/k}(0)$ and $y$. We are trying to show that the Gromov-Hausdorff limit is $\mathbb{E}^{N}$, which amounts to showing that
$\lim_{k\rightarrow\infty}d_{GH}\big{(}\big{(}\overline{B}_{r}(x_{k}),d_{X_{k}}\big{)},\big{(}\overline{B}_{r}(0),d_{E}\big{)}\big{)}=0,$
where $d_{E}$ is the Euclidean distance function. Actually, the notion of pointed Gromov-Hausdorff convergence is not necessary here because of Lemma 7.7 of [2], which states that for an $N$-dimensional pulled thread space $X$ where the thread is a ball of diameter $D$
$d_{GH}((X,d_{X}),(\mathbb{E}^{N},d_{E}))\leq 2D.$
(The lemma is proved by finding a suitable correspondence). However, in the limit the diameter of the ball goes to zero, so we have
$\lim_{k\rightarrow\infty}d_{GH}((X_{k},d_{X_{k}}),(\mathbb{E}^{N},d_{E}))=0$
as required. In general, the Gromov-Hausdorff limit $X_{\infty}$ will not necessarily be $\mathbb{E}^{N}$ and there will be a possibly countable collection of disjoint compact connected sets in each smocking set in the sequence, so it was necessary to show that a formal argument with correspondences goes through in the general case.
It is possible to think of potential counterexamples which are something as follows: one imagines a smocking pattern given by alternating balls of radius $1/k$ with balls of radius $(2^{k}-2)/2^{k}$. In the case of $k=1$, this implies that $S_{1}$ is a pattern which alternates between unit $N$-balls and points. In the limit as $k$ goes to infinity, $S_{k}$ will become the opposite pattern, since the unit balls in $S_{1}$ will all shrink to points and the points will expand out to be unit balls. This will give Hausdorff convergence to a smocking set $S_{\infty}$ which can be smocked to obtain a well-defined smocked space $X_{\infty}$.
However, $(2^{k}-2)/2^{k}$ approaches $1$ faster than $1/k$ approaches zero as $k$ goes to infinity. It follows that if the balls in $S_{1}$ are sufficiently close to the points, at some stage in the sequence of smocking sets $S_{k}$ the growing ball will merge with the shrinking ball to form a new set which might develop such that the difference between the distance between two points in $X_{k}$ and the distance between the corresponding two points in $X_{k+1}$ at some stage momentarily gets larger, whereas one would expect the difference between distances between pairs of points to decrease with increasing $k$ if there is to be Gromov-Hausdorff convergence of the metric spaces as $k$ tends to infinity. This fails to be a counterexample because by hypothesis the family of subsets over $\mathbb{E}^{N}$ need to remain disjoint from each other as the sequence $\{S_{k}\}_{k=1}^{\infty}$ converges to its Hausdorff limit.
3.2 Subsequential Gromov-Hausdorff convergence
The next question regarding Gromov-Hausdorff convergence is when a sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ has a Gromov-Hausdorff converging subsequence. Gromov’s pre-compactness theorem gives us sufficient conditions for a sequence of metric spaces to have such a subsequence, but the theorem requires a uniform upper bound on diameter and the number of disjoint balls of any given radius (intuitively, one should be able to fill the space with not too many small disjoint balls as one goes to smaller and smaller scales) [6]. Smocked spaces have unbounded diameter and do not satisfy these hypotheses, and so it is not clear if there should be a result on subsequential Gromov-Hausdorff convergence for these spaces.
In [2], it was shown that sequences of re-scalings of smocked spaces do converge in the pointed Gromov-Hausdorff sense .even though the spaces themselves do not satisfy the hypotheses of the Gromov pre-compactness theorem. There are some immediate examples of Gromov-Hausdorff converging subsequences which one can construct. For example, take a sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ such that $X_{k}$ is a smocked space with one single smocking stitch for odd $k$ and such that $X_{k}$ is a smocked space with two smocking stitches which are the same compact subsets $\mathbb{E}^{n}$ for each even $k$. This sequence of smocked spaces contains a Gromov-Hausdorff converging subsequence given by $(X_{2k})$.
Generally speaking, we would like conditions for a sequence of smocked spaces to contain a pointed Gromov-Hausdorff converging subsequence. It turns out that to get such conditions it is sufficient to assume uniform bounds on the smocking constants. Intuitively, the result says that one cannot pack in too many sets too closely together in the smocked space. This may explain why the examples found by Sormani et al. can be shown to have rescaled sequences which converge even though smocked spaces themselves do not meet the hypotheses of Gromov’s compactness theorem. As stated earlier, one obtains a sequence of metric spaces when a metric space is repeatedly rescaled and the Gromov-Hausdorff limit (if it exists) is the tangent cone at infinity.
Proposition 3.2: A sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ contains a Gromov-Hausdorff converging subsequence if there are uniform upper bounds on the smocking constants.
Proof: The proof is similar to others in the literature, but we will reproduce the details here for completeness [1]. We must prove that there is a finite number of indices $i_{1},...,i_{p}$ such that for every $k$, there exists $j$ with
$d_{GH}((B_{r}(x_{0}^{(k)}),dX_{k}),(B_{r}(x_{0}^{(i_{j})}),dX_{i_{j}}))\leq\text{constant}\cdot D_{k},$
where $D_{k}$ is defined to be $\text{sup}\{L_{\text{min}}^{k},L_{\text{max}}^{k},h_{k},\delta_{k}\}$. $D_{k}$ is assumed to be finite by hypothesis.
Since $D_{k}$ is finite, for each ball $(B_{r}(x_{0}^{(k)}),dX_{k})$ the maximal disjoint filling is achieved by balls of radius $cD_{k}$ centred at points of a $2cD_{k}$-net of the ball (call the net $L$), for some $c>0$. The number of balls in the filling is bounded by some quantity $N=N(D_{k},x_{0}^{(k)},r)$. A choice of bijection $f:\{1,...,N\}\longrightarrow L$ induces a distance function $d_{L}$ on the set $\{1,...,N\}$. The function $d_{L}:\{1,...,N\}\times\{1,...,N\}\longrightarrow\mathbb{R}_{+}$ which assumes values between $2cD_{k}$ and $2r$ on the subset of $S$ of pairs of different integers has $N^{2}-N$ elements. The space $[2cD_{k},2r]^{N^{2}-N}$ is precompact with respect to the product distance function, so there exist finitely many indices $i_{1},...,i_{p}$ such that, for each $i$, there exists $j$ such that
$\sup_{S}|d_{k}(m,q)-d_{i_{j}}(m,q)|\leq c^{2}D_{k}^{2},$
which implies that
$\sup_{S}\frac{d_{k}(m,q)}{d_{i_{j}}(m,q)}\leq\frac{1+cD_{k}}{1-cD_{k}}.$
This in turn implies that
$d_{H}((B_{r}(x_{0}^{(k)}),dX_{k}),(B_{r}(x_{0}^{(i_{j})}),dX_{i_{j}}))\leq(2r+4)cD_{k}.$
Finally, Hausdorff convergence here implies pointed Gromov-Hausdorff convergence, so we have
$d_{GH}((B_{r}(x_{0}^{(k)}),dX_{k}),(B_{r}(x_{0}^{(i_{j})}),dX_{i_{j}}))\leq(2r+4)cD_{k},$
[…]
Remark 3.3: Note that having uniform bounds on the smocking constants is a sufficient, not a necessary, condition for subsequential convergence. To show that it is not necessary, one may consider a sequence where each space $X_{k}$ consists of two ’stitches’ which move further apart with increasing $k$. This converges to $\mathbb{E}^{N}$ in the limit, but the smocking separation factor $\delta$ is unbounded.
3.3 Normed spaces from smocked spaces
Finally, we show that one can always obtain a finite-dimensional normed vector space as a pointed Gromov-Hausdorff limit of a sequence of smocked spaces. In general, there are examples of smocked spaces which do not have unique tangent cones at infinity, nor is the cone always a normed space [2].
Proposition 3.4: Given a finite-dimensional normed vector space $(Y,d_{Y})$, there exists a smocked space $(X,d_{X})$ whose tangent cone at infinity is $(Y,d_{Y})$.
Proof: For a metric space $(X,d)$ to be a normed space, the metric $d$ must be translation invariant
$d(x,y)=d(x+a,y+a),\>\>\>\>\>\>\>\forall a\in X$
and homogeneous
$d(cx,cy)=cd(x,y),\>\>\>\>\>\>\>\forall c\in\mathbb{R}.$
[…]
4 Smocked spaces with measures
4.1 Measured convergence of smocked measure spaces
A natural question is whether the results which were derived above still apply for metric spaces equipped with measures. We will begin by recalling the definition of metric measure spaces given in [7].
Definition 4.1 A metric measure space is a triple $(X,d,\mu)$, where $(X,d)$ is a metric space and $\mu$ is a measure on the Borel $\sigma$-algebra of $X$.
We then propose to extend this definition to smocked spaces in order to work with what we will call pointed smocked metric measure spaces (or smocked measure spaces for short). As before, we must use pointed convergence because smocked spaces are non-compact.
Definition 4.2 A smocked metric measure space is a triple $(X,d,\mu)$, where $(X,d)$ is a smocked space and $\mu$ is a non-zero Borel measure which is finite on bounded sets. A pointed smocked metric measure space $(X,d,\mu,\overline{x})$ is a smocked metric measured space together with a given point $\overline{x}\in\text{supp}(\mu)\subset X$.
The first question we might ask is whether we can obtain analogues of our previous convergence results established above, but now for the case of smocked measure spaces. We will begin by defining a suitable notion of pointed measured Gromov-Hausdorff convergence used by Lott and Villani and adapted by Gigli, Mondino and Savaré [8-10]. This notion works with proper (that is, bounded closed subsets are compact) pointed metric measure spaces. $\mathbb{E}^{N}$ is proper finite-dimensional, so distances converge uniformly on bounded sets.
Definition 4.3 If one has a sequence of pointed smocked metric measure spaces $(X_{j},d_{j},\mu_{j},\overline{x}_{j})$, pointed measured Gromov-Hausdorff convergence is defined as follows:
$(X_{j},d_{j},\mu_{j},\overline{x}_{j})\xrightarrow{pmGH}(X_{\infty},d_{\infty},\mu_{\infty},\overline{x}_{\infty})$
if and only if for every $\epsilon,R>0$ there exists $N(\epsilon,R)$ such that, for all $n\geq N$, there exists a Borel map $f_{j}^{R,\epsilon}$: $B_{R}(\overline{x}_{j})\rightarrow\overline{x}_{\infty}$ such that
(i) $f_{j}^{R,\epsilon}(\overline{x}_{j})=\overline{x}_{\infty}$;
(ii) $\text{sup}_{x,y\in B_{R}(\overline{x}_{j})}|d_{j}(x,y)-d_{\infty}(f_{j}^{R,\epsilon}(x),(f_{j}^{R,\epsilon}(y))|\leq\epsilon$;
(iii) the $\epsilon$-neighbourhood of $f_{j}^{R,\epsilon}(B_{R}(\overline{x}_{j}))$ contains $B_{R-\epsilon}(\overline{x}_{\infty})$;
(iv) $(f_{j}^{R,\epsilon})_{\#}(\mu_{j}\rvert_{B_{R}(\overline{x}_{j})})$ weakly converges to $\mu_{\infty}\rvert_{B_{R}(\overline{x}_{\infty})}$ as $n\rightarrow\infty$, for almost every $R>0$.
This definition is somewhat more complicated than the one for pointed Gromov-Hausdorff convergence, but we can still show that if one has a sequence of smocking sets $S_{k}$ which Hausdorff converge to $S_{\infty}$ as before, then we can always define corresponding smocked measure spaces $(X_{k},\mu_{K})$ which converge to a pointed measured Gromov-Hausdorff limit. By Proposition 3.1, it is only necessary to show that pointed Gromov-Hausdorff convergence implies pointed measured Gromov-Hausdorff convergence.
Proposition 4.4: If a sequence of smocked spaces $(X_{k},x_{k},d_{k})$ converges in the pointed Gromov-Hausdorff sense to $(X_{\infty},x_{\infty},d_{\infty})$, then one can define a sequence of smocked measure spaces $(X_{k},d_{k},\mu_{k},\overline{x}_{k})$ which converges in the pointed measured Gromov-Hausdorff sense to $(X_{\infty},d_{\infty},\mu_{\infty},\overline{x}_{\infty})$.
[…]
The following corollary is immediate.
Corollary 4.5: If one defines a sequence of smocked measure spaces using smocking sets $S_{k}\subset\mathbb{E}^{N}$ such that the sets Hausdorff converge in the limit to a smocking set $S_{\infty}$, then one can define corresponding smocked measure spaces which converge to the corresponding limit space in the pointed measured Gromov-Hausdorff sense.
We conjecture that it should be possible to show sufficient conditions for subsequential pointed measured Gromov-Hausdorff convergence of smocked measure spaces by explicitly constructing an abstract limit space with an ancillary measure and showing that convergence is obtained after extraction of subsequences.
Conjecture 4.6: A sequence of smocked measure spaces $\{(X_{k},d_{k},\mu_{k},\overline{x}_{k})\}_{k=1}^{\infty}$ contains a subsequence which converges in the pointed measured Gromov-Hausdorff sense if there are uniform upper bounds on both the smocking constants and $\text{sup}\>\mu_{k}(B_{r}(x_{0}^{(k)}))$.
Remark 4.7: Much recent work has focussed on stability of metric measure spaces under different types of convergence. As an example, one may consider stability of the curvature dimension condition under pointed measured Gromov-Hausdorff convergence [13, 14, 15-17,10]. There is no such stability result for smocked measure spaces because they are not length spaces nor are they compact, which means that the associated Wasserstein spaces also cannot be compact or length spaces.
4.2 Smocked Wasserstein Spaces
Pulled thread spaces (and more generally, smocked spaces) can be defined by starting with any geodesic metric space or length space [1]. We conjecture here that Wasserstein spaces which are also length spaces can be smocked by considering smocking sets which are disjoint connected compact subset of the space of Borel probability measures. It is possible that smocking the Wasserstein spaces would have an influence on $GW_{p}$-convergence of the original metric measure spaces.
Another basic principle of optimal transport is that a pair of Borel probability measures on a manifold which evolve under a diffusion equation should converge in the Wasserstein distance if the manifold satisfies a curvature condition (usually positive Ricci curvature as defined above) [19]. The work of Sturm and von Renesse shows that if a smooth, compact Riemannian manifold $M$ is equipped with two Borel probability measures $\mu$ and $\nu$ which spread according to the heat equation, then the lower bound on the Ricci curvature of $M$ controls the speed with which the Wasserstein distance between the two measures can expand or decay over time. In particular, it was shown that the following properties are all equivalent for such a manifold:
Proposition 4.8 [20]: For $K\in\mathbb{R}$ and a compact Riemannian manifold $M$, the following are equivalent:
(i) $\text{Ric}(M)\geq K$.
(ii) For all $x$ and $y$ in $M$ and all $t>0$, there exists $r\in[1,\infty]$ with
$d_{W_{P}}(p_{t}(x,\cdot),p_{t}(y,\cdot))\leq e^{-Kt}\cdot d(x,y).$
(iii) For all $r\in[1.\infty]$, all $\mu$ and $\nu$ in $\mathcal{P}_{P}(M)$, and all $t>0$
$d_{W_{P}}(\mu p_{t}(x,\cdot),\nu p_{t}(y,\cdot))\leq e^{-Kt}\cdot d_{W_{P}}(\mu,\nu),$
where $p_{t}(x,y)$ denotes the minimal positive fundamental solution of the heat equation.
We conjecture that if a smocked Wasserstein distance can be defined, condition (iii) can be replaced with the same statement but using this new distance.