### v3

# Gromov–Hausdorff convergence and tangent cones of smocked spaces

Abstract
Smocked spaces are a class of metric spaces which were introduced to generalize pulled thread spaces. We investigate convergence of these spaces, showing that if the underlying smocking sets converge in Hausdorff distance and satisfy local uniform bounds on the smocking constants, then the associated smocked spaces converge in the pointed Gromov-Hausdorff sense. We prove a corresponding precompactness result using a similar assumption on the smocking constants. We also show that every finite-dimensional normed vector space arises as the tangent cone at infinity of a suitably constructed smocked space. Finally, we extend the convergence theory to the setting of smocked metric measure spaces, proving stability under pointed measured Gromov-Hausdorff convergence. These results establish a basic geometric framework for smocked spaces and demonstrate that they exhibit controlled limit behavior.
Keywords: Smocked spaces, Gromov–Hausdorff convergence, metric geometry, Hausdorff convergence, tangent cones, metric measure spaces.
53C23, 54E35; Secondary 28A33, 51F30.
1 Introduction
The concept of a pulled thread space has existed for some time in metric geometry [1]. Roughly speaking, the idea is to take Euclidean space $\mathbb{E}^{N}$, view the space as a piece of cloth, sew along an interval or some other compact connected set in that space, and then pull the thread tight to obtain a new metric space where the sewing interval is now viewed as a point (a formal definition will be given shortly). Building on this concept, Sormani et al. recently introduced the notion of a smocked metric space and studied the balls and geodesics for several such spaces [2]. The definition is lengthy and will be given in Section 2, but one may think of a smocked space as a generalisation of a pulled thread space to include a (possibly countable) collection of disjoint compact connected sets on $\mathbb{E}^{N}$, all of which are ’pulled tight’.
One can prove that these spaces are well-defined as metric spaces and meet all the relevant axioms, but there remains a number of basic questions about convergence of these metric spaces and their subsequences, which we will address in this article [2]. One of the immediate applications mentioned in [2] is to apply results on Sormani-Wenger intrinsic flat convergence of smocked spaces to an open question of Gromov and Sormani on tori of almost nonnegative scalar curvature [3, 4]. In this article, we will not consider intrinsic flat convergence and will instead prove some rigorous results only for Gromov-Hausdorff convergence. It would certainly be desirable to have some more results of this kind before embarking on further studies of smocked spaces and their tangent cones at infinity [5]. Although a variety of smocked spaces and their tangent cones were studied in [2], it was necessary to carefully consider the explicit formulae for the distance functions in each case. In this article, we wish to show that it is possible to do analysis and geometry in smocked spaces in a general sense without having to know the distance function a priori.
The article is structured as follows. In Section 2, we state basic definitions and results. In Section 3, we give a counterexample which shows that Hausdorff convergence of a sequence of smocking sets $S_{k}$ to a limit set $S_{\infty}$ does not necessarily imply that the corresponding smocked spaces $X_{k}$ converge to $X_{\infty}$ in the pointed Gromov-Hausdorff sense, answering an open question of Sormani et al. in the negative. We show that the smocked spaces do converge if one assumes local Hausdorff convergence on balls and local control of the smocking constants used to define a smocking space. We show that a similar story also holds for subsequential Gromov-Hausdorff convergence: global bounds on the smocking constants are sufficient for precompactness, but one must assume stronger local uniform bounds in order to have necessity. We end the section by showing that for any given finite-dimensional normed space, there is always a smocked space whose unique tangent cone at infinity is that space. In Section 4, we show that these convergence results extend naturally to smocked spaces equipped with measures. We finish with conclusions and future directions in Section 5.
2 Preliminaries
2.1 Gromov-Hausdorff convergence
In this subsection, we present basic definitions concerning limits with the Gromov-Hausdorff distance.
Definition 2.1: Given two metric spaces $X$ and $Y$, the Gromov-Hausdorff distance is defined to be
$d_{GH}(X,Y)=\text{inf}\bigg\{d_{H}^{Z}(f(X),f(Y))\bigg\}$
taken over all metric spaces $Z$ and isometric embeddings $f:X\rightarrow Z$ and $g:Y\rightarrow Z$. $d_{H}^{Z}$ is the Hausdorff distance between subsets of $Z$
$d_{H}^{Z}(A,B)=\text{inf}\{\epsilon>0:B\subset T_{\epsilon}(A)\>\text{and}\>A\subset T_{\epsilon}(B)\},$
where $T_{\epsilon}(A)=\{x\in Z:d_{Z}(x,A)<\epsilon\}$.
Definition 2.2: If one has a sequence of complete noncompact metric spaces $(X_{j},d_{j})$ and points $x_{j}\in X_{j}$, pointed Gromov-Hausdorff convergence is defined as follows:
$(X_{j},d_{j},x_{j})\xrightarrow{pGH}(X_{\infty},d_{\infty},x_{\infty})$
if and only if for every $R>0$ the closures of the balls of radius $R$ in $X_{j}$ converge in the Gromov-Hausdorff sense as metric spaces with the restriction of the distance function to closures of balls in $X_{\infty}$
$(\overline{B}_{R}(x_{j}),d_{j})\xrightarrow{GH}(\overline{B}_{R}(x_{\infty}),d_{\infty}).$
A sequence of metric spaces can be obtained by repeatedly rescaling a space. Although some spaces like the taxi space are isometric to their rescalings, this is generally not the case. If a sequence or subsequence obtained by rescaling converges in the Gromov-Hausdorff sense, the limit space obtained is known as the tangent cone at infinity.
Definition 2.3: A complete non-compact metric space $(X,d_{X})$ with infinite diameter has a tangent cone at infinity $(Y,d_{Y})$ if there is a sequence of rescalings $R_{j}\rightarrow\infty$ and points $x_{0}\in X$ and $y_{0}\in Y$ such that
$(X,d/r_{j},x_{0})\xrightarrow{pGH}(Y,d_{Y},y_{0}).$
2.2 Smocked spaces
This section contains definitions and main results regarding smocked spaces.
Definition 2.4 Start with a finite or countable collection of disjoint connected compact sets in $\mathbb{E}^{N}$ called smocking stitches
$I=\{I_{j}:j\in J\}.$
Also, assume the existence of a positive smocking separation factor $\delta$
$\delta=\text{min}\{|z-z^{\prime}|:z\in I_{j},z^{\prime}\in I_{j^{\prime}},j\neq j^{\prime}\in J\}>0.$
Using these one can define a smocked metric space $(X,d)$ where each stitch in the collection is viewed as a single point:
[…]
where the smocking set $S$ is
$S=\bigcup_{j\in J}I_{j}.$
We also have a smocking map $\pi:\mathbb{E}^{N}\rightarrow X$ defined by
[…]
for some index set $J$ and a smocked distance function $d:X\times X\rightarrow[0,\infty)$ which is defined for $x,y\notin\pi(S)$ and stitches $I_{m}$ and $I_{n}$
$d(x,y)=\text{min}\{d_{0}(x,y),d_{1}(x,y),$...\},
$d(x,I_{n})=\text{min}\{d_{0}(x,z),d_{1}(x,z),...:z\in I_{n}\},$
$d(I_{m},I_{n})=\text{min}\{d_{0}(z^{\prime},z),d_{1}(z^{\prime},z),...:z^{\prime}\in I_{m},z\in I_{n}\},$
where $d_{k}$ is the sum of lengths of segments which go between $k$ stitches:
$d_{k}(v,w)=\text{min}\{|v-z_{1}|+\sum_{i=1}^{k-1}|z^{\prime}_{i}-z_{i+1}|+|z^{\prime}_{k}-w|:z_{i},z^{\prime}_{i}\in I_{j_{i}},j_{1}\neq...\neq j_{k}\in J\},$
for any $v,w\in\mathbb{E}^{n}$. The smocking pseudometric $\overline{d}:\mathbb{E}^{N}\times\mathbb{E}^{N}\rightarrow[0,\infty)$ is defined as
$\overline{d}(v,w)=d(\pi(v),\pi(w))=\text{min}\{d_{k}(v^{\prime},w^{\prime}):\pi(v)=\pi(v^{\prime}),\pi(w)=\pi(w^{\prime}),k\in\mathbb{N}\}.$
The definition of a smocked space refers to a set of four constants, which are referred to in [2] as the smocking constants: these are the smocking depth $h$, the maximum smocking length $L_{\text{max}}$, the minimum smocking length $L_{\text{min}}$, and the smocking separation factor $\delta$, which was already defined in Definition 2.6.
Definition 2.5 The smocking depth $h$ is defined as
$h=\text{inf}\{r:\forall x\in X\>\exists j\in J\>\exists z\in I_{j}\>\text{such that}\>|x-z|<r\}.$
Definition 2.6 The smocking lengths $L_{\text{max}}$ and $L_{\text{min}}$ are defined using the diameters of the smocking stitches as follows:
$L_{\text{min}}=\text{inf}\{D(I_{j}):\>j\in J\},$
$L_{\text{max}}=\text{sup}\{D(I_{j}):\>j\in J\}.$
The main theorem proved in [2] for smocked spaces is the following result:
Proposition 2.7: Given an $N$-dimensional smocked space $(X,d)$ such that
$|\overline{d}(x,x^{\prime})-|F(x)-F(x^{\prime})||\leq K,\>\>\>\>\>\>\forall\>x,x^{\prime}\in\mathbb{E}^{N},$
where $F:\mathbb{E}^{N}\rightarrow[0,\infty)$ is a norm, then $(X,d)$ has a unique tangent cone at infinity $(\mathbb{R}^{N},d_{F})$, where
[…]
2.3 Metric geometry
The proofs of the results in Section 3 will require the definition of a correspondence for two metric spaces $X$ and $Y$:
Definition 2.8: Given two metric spaces $X$ and $Y$, a correspondence $C$ is a set $C\subset X\times Y$ such that
$\forall x\in X\>\exists y\in Y\>\text{such that}\>(x,y)\in C\>\text{and}\>\forall y\in Y\>\exists x\in X\>\text{such that}\>(x,y)\in C.$
Some of the proofs will also require the definition of an $\epsilon$-net.
Definition 2.9 An $\epsilon$-net in a pointed metric space $(X,x)$ is a subset $L$ of $X$ containing $x$ whose $\epsilon$-neighbourhood covers $X$ and for which there is some $\epsilon^{\prime}>0$ with $d(l_{1},l_{2})\geq\epsilon^{\prime}$ for all $l_{1}\neq l_{2}$ in $L$.
We will also use two additional elementary lemmas which we prove here (these are used implicitly in the proof of Proposition 3.3).
Lemma 2.10 Let $\{S_{\alpha}\}$ be a collection of pairwise disjoint compact subsets (stitches) of $\overline{B_{R+1}(0)}\subset\mathbb{E}^{N}$ such that each $S_{\alpha}$ has diameter $\leq L$ and distinct stitches are separated by distance at least $\delta_{0}>0$. Then any rectifiable curve $\tilde{\gamma}\subset\overline{B_{R+1}(0)}$ of Euclidean length $\ell(\tilde{\gamma})\leq L_{0}$ meets at most
$M\;:=\;1+\Big\lfloor\frac{L_{0}}{\delta_{0}}\Big\rfloor$
distinct stitches. In particular, for $L_{0}=2R+2$ the number of stitches intersected is bounded by a constant $M(R,L,\delta_{0})$ depending only on $R,L,\delta_{0}$.
Proof: Order the distinct stitches met by $\tilde{\gamma}$ in the sequence encountered along the curve $S_{\alpha_{1}},S_{\alpha_{2}},\dots,S_{\alpha_{m}}$. For each consecutive pair $S_{\alpha_{j}},S_{\alpha_{j+1}}$ choose points $p_{j}\in S_{\alpha_{j}}$ and $q_{j}\in S_{\alpha_{j+1}}$ which lie along the curve in that order. By the separation hypothesis, we assume that $|p_{j}-q_{j}|_{\mathbb{E}^{N}}\geq\delta_{0}$. Summing over the m1m-1 gaps, we obtain
$\ell(\tilde{\gamma})\ \geq\ \sum_{j=1}^{m-1}|p_{j}-q_{j}|\ \geq\ (m-1)\delta_{0}.$
[…]
Lemma 2.11 Let $(X,d_{X})$ be a smocked space obtained from $\mathbb{E}^{N}$ by collapsing a collection of compact stitches $\{S_{\alpha}\}$, and let $\pi:\mathbb{E}^{N}\to X$ be the quotient map. Assume that in $\overline{B_{R+1}(0)}$ every stitch has diameter $\leq L$ and distinct stitches are $\delta_{0}$–separated. Then for every $r>0$ there exists $R(r)<\infty$, depending only on $r,L,\delta_{0}$ and $R$, such that
$\pi^{-1}\big(B_{r}^{X}(\pi(0))\big)\subset B_{R(r)}^{\mathbb{E}^{N}}(0).$
In particular, when we assume uniform local bounds on the smocking constants for the sequence $\{X_{k}\}$, this $R(r)$ can be chosen uniformly in $k$.
[…]
3 Gromov-Hausdorff convergence of smocked spaces
In this section, we will investigate general results on Gromov-Hausdorff convergence, including sufficient and necessary conditions for precompactness and whether or not Hausdorff convergence of the smocking sets implies Gromov-Hausdorff convergence of the smocked spaces. We also prove a general result for tangent cones at infinity. The theory developed in this section thus answers three open questions of Sormani et al.
3.1 Relation between Hausdorff and Gromov-Hausdorff convergence
An open question for these spaces is whether Hausdorff convergence of the smocking sets $S_{k}$ to $S_{\infty}$ by itself implies Gromov–Hausdorff convergence of the associated smocked spaces $X_{k}$ to $X_{\infty}$ [2]. This is false in general. The key point is that Hausdorff convergence does not by itself control the amount of length that can be collapsed and hidden inside a stitch, which influences distances in the smocked space. This can be demonstrated with an explicit counterexample (the example is one-dimensional, but the construction can easily be generalized).
Example 3.1: Fix the ambient ball radius $R=1$ and the ambient interval $I=[-1,1]\subset\mathbb{R}$. We will construct a sequence of families of disjoint closed intervals $\{S_{\alpha}^{(k)}\}\subset I$ so that the unions $U_{k}:=\bigcup_{\alpha}S_{\alpha}^{(k)}$ converge in the Hausdorff distance to $U:=I$, but the associated smocked spaces $X_{k}$ do not admit a subsequence which converges in the pointed Gromov-Hausdorff sense.
Fix two distinct numbers $L^{\mathrm{odd}},L^{\mathrm{even}}\in(0,2)$. For example, we may use $L^{\mathrm{even}}=1/3$ and $L^{\mathrm{odd}}=2/3$. For each integer $k\geq 1$, set
$N_{k}:=k,\qquad\delta_{k}:=\frac{1}{N_{k}+1}.$
We build $N_{k}$ disjoint small intervals in $I$ so that the maximum length of a gap between adjacent intervals is less than or equal to $\delta_{k}$, hence the Hausdorff distance from $U_{k}$ to $I$ will be $\leq\delta_{k}$ and tends to 0 as $k\to\infty$.
Choose a sequence of total collapsed lengths $L_{k}$ which alternates as follows:
$L_{k}=\begin{cases}L^{\mathrm{even}},&k\ \text{even},\\[4.0pt] L^{\mathrm{odd}},&k\ \text{odd}.\end{cases}$
For each $k$, we distribute the total length $L_{k}$ amongst the $N_{k}$ small intervals approximately equally: set each stitch length $\ell_{j,k}=L_{k}/N_{k}$ for $j=1,\dots,N_{k}$. Since $\ell_{j,k}=L_{k}/N_{k}\to 0$ as $k\to\infty$, the individual stitch diameters tend to 0.
Next, place the $N_{k}$ intervals in $I$ in increasing order so that the gaps between them are all $\leq\delta_{k}$ (this is possible because $N_{k}$ is finite and we may choose the starting position as we wish). Concretely, take the leftmost interval to begin at 1-1 and then place intervals and small gaps alternately so that the number of gaps is $N_{k}+1$ and each gap has length at most $\delta_{k}$. This is possible because
$(N_{k}+1)\delta_{k}=(N_{k}+1)\frac{1}{N_{k}+1}=1,$
so the total length of gaps can be arranged to be less than or equal to 1, leaving enough room for the total stitch length $L_{k}\leq 2$. With minor endpoint adjustments, the construction fits in $I$ for all large $k$. By construction, the maximal gap size in $I\setminus U_{k}$ is $\leq\delta_{k}$, hence $d_{H}(U_{k},I)\leq\delta_{k}\to 0$, so $U_{k}$ converges to $I$ in the Hausdorff metric on compact subsets.
Let $\pi_{k}:\mathbb{R}\to X_{k}$ denote the smocking map which collapse each stitch interval $S_{k}$ to a single point. For any two points, $a<b\in\mathbb{R}$ the smocked space distance $d_{X_{k}}(\pi_{k}(a),\pi_{k}(b))$ is the infimum length of paths joining $a$ to $b$ in $\mathbb{R}$ measured with the convention that any portion lying inside a collapsed stitch contributes zero length. In particular, for the endpoints $1,1\in I$ we have the exact formula
$d_{X_{k}}\big(\pi_{k}(-1),\pi_{k}(1)\big)\;=\;2-L_{k}.$
The interval [1,1][-1,1] has length $2$, and collapsing the union of disjoint subintervals of total length $L_{k}$ reduces the distance by exactly $L_{k}$.
However, since $L_{k}$ alternates between the two distinct values $L^{\mathrm{even}}$ and $L^{\mathrm{odd}}$, the sequence of real numbers given by $d_{X_{k}}$ does not converge (it attains two distinct limit points for the even and the odd subsequences), hence the pointed metric spaces $(X_{k},\pi_{k}(0))$ do not converge to $(X_{\infty},\pi_{\infty}(0))$ in the pointed Gromov–Hausdorff topology.
If we rule out pathological counterexamples, then generally we would expect Hausdorff convergence of the smocking sets to imply Gromov-Hausdorff convergence of the corresponding smocked spaces. A well-behaved example where this is true is provided below.
Example 3.2: Start with $\mathbb{E}^{N}$ and take the sequence of smocking sets $S_{k}$ to be the sequence of closed balls about the origin with radius $1/k$. In the limit, $S_{\infty}$ is the point at the origin, so $X_{\infty}=\mathbb{E}^{N}$, since we have removed a point and replaced it with a point. Now for the sequence of smocked spaces, consider $X_{k}$:
[…]
In this space $\overline{B}_{1/k}(0)$ is identified with a point, so the distance between two points $x$ and $y$ is the minimum of the direct distance between $x$ and $y$ and all the possible distances between $x$ and $z\in\overline{B}_{1/k}(0)$ plus the distances between $z^{\prime}\in\overline{B}_{1/k}(0)$ and $y$. We are trying to show that the Gromov-Hausdorff limit is $\mathbb{E}^{N}$, which amounts to showing that
$\lim_{k\rightarrow\infty}d_{GH}\big(\big(\overline{B}_{r}(x_{k}),d_{X_{k}}\big),\big(\overline{B}_{r}(0),d_{E}\big)\big)=0,$
where $d_{E}$ is the Euclidean distance function. Actually, the notion of pointed Gromov-Hausdorff convergence is not necessary here because of Lemma 7.7 of [2], which states that for an $N$-dimensional pulled thread space $X$ where the thread is a ball of diameter $D$
$d_{GH}((X,d_{X}),(\mathbb{E}^{N},d_{E}))\leq 2D.$
(The lemma is proved by finding a suitable correspondence). However, in the limit the diameter of the ball goes to zero, so we have
$\lim_{k\rightarrow\infty}d_{GH}((X_{k},d_{X_{k}}),(\mathbb{E}^{N},d_{E}))=0$
as required.
Although Hausdorff convergence of the smocking sets does not generally imply Gromov-Hausdorff convergence of the smocking spaces, we can instead prove a natural local replacement: Hausdorff convergence for each bounded ball along with a uniform local bound on the smocking constants is sufficient to imply pointed Gromov-Hausdorff convergence.
Proposition 3.3: Let $\{S_{\alpha}^{(k)}\}_{\alpha\in\mathcal{A}_{k}}$ be the collections of compact stitches in $\mathbb{E}^{N}$ defining smocked spaces $X_{k}=\mathbb{E}^{N}/\!\sim_{k}$ with smocking maps $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ and basepoints $x_{k}=\pi_{k}(0)$. Assume:
  1. (H1)
    (Local Hausdorff convergence) For every $R>0$, the Hausdorff distance between unions of stitches meeting $\overline{B_{R}(0)}$ converges to zero:
    $d_{H}\Big(\bigcup_{\alpha:S_{\alpha}^{(k)}\cap\overline{B_{R}}\neq\varnothing}S_{\alpha}^{(k)},\ \bigcup_{\beta:S_{\beta}^{(\infty)}\cap\overline{B_{R}}\neq\varnothing}S_{\beta}^{(\infty)}\Big)\xrightarrow[k\to\infty]{}0.$
  2. (H2)
    (Local uniform bounds on smocking constants) For every $R>0$, there exist $L_{R}<\infty$, $\delta_{R}>0$ and $K_{R}$ such that for all $k\geq K_{R}$ every stitch intersecting $\overline{B_{R}(0)}$ has diameter $\leq L_{R}$ and distinct such stitches are separated by Euclidean distance at least $\delta_{R}$.
Let $X_{\infty}$ be the smocked space obtained from the limiting stitches $\{S_{\beta}^{(\infty)}\}$, with basepoint $x_{\infty}=\pi_{\infty}(0)$. Then $(X_{k},x_{k})\xrightarrow{\mathrm{pGH}}(X_{\infty},x_{\infty})$ as $k\to\infty$.
Proof: Fix $R>0$ and $\varepsilon>0$. The strategy of the proof is to produce a correspondence between $B_{R}^{X_{k}}(x_{k})$ and $B_{R}^{X_{\infty}}(x_{\infty})$ of distortion $\varepsilon$, for all $k$. By (H1) and (H2) we can choose $\eta>0$ and $K$ so that for $k\geq K$ the finite collections of stitches meeting $\overline{B_{R+1}(0)}$ are in bijection (matched pairwise) and each matched pair is Hausdorff-close by $\eta$. Furthermore, each such stitch has diameter $\leq L:=L_{R+1}$ and distinct stitches are $\geq\delta_{0}:=\delta_{R+1}$ apart. Let $N$ be a uniform bound on the number of stitches meeting $\overline{B_{R+1}(0)}$. This bound exists by (H2).
Define a correspondence $\mathcal{R}\subset X_{k}\times X_{\infty}$ by declaring that $(x_{k},x_{\infty})\in\mathcal{R}$ if and only if one of the following holds:
  1. 1.
    there exists $z\in\overline{B_{R+1}(0)}$ such that $x_{k}=\pi_{k}(z)$ and $x_{\infty}=\pi_{\infty}(z)$; or
  2. 2.
    there is a matched pair of stitches $S_{\alpha}^{(k)}\leftrightarrow S_{\beta}^{(\infty)}$ such that $x_{k}=\pi_{k}(z_{k})$, $x_{\infty}=\pi_{\infty}(z_{\infty})$ for some $z_{k}\in S_{\alpha}^{(k)}$, $z_{\infty}\in S_{\beta}^{(\infty)}$.
The idea of the construction is to pair up points in $X_{k}$ and $X_{\infty}$ that are images of the same preimage $z\in\mathbb{E}^{N}$, and to pair each stitch in $X_{k}$ with its matched limiting stitch in $X_{\infty}$. The projections of $\mathcal{R}$ onto each factor are surjective onto the images of $\overline{B_{R+1}(0)}$ under $\pi_{k}$ and $\pi_{\infty}$, respectively. In particular, $\mathcal{R}$ induces a correspondence between $B_{R}^{X_{k}}(x_{k})$ and $B_{R}^{X_{\infty}}(x_{\infty})$. The margin of +1+1 ensures that every point of the metric balls admits a lift within $\overline{B_{R+1}(0)}$.
Let $(u_{i},v_{i})\in\mathcal{R}$ with lifts $a_{i}\in\overline{B_{R+1}(0)}$ so that $u_{i}=\pi_{k}(a_{i})$ and $v_{i}=\pi_{\infty}(a_{i})$, except for when $a_{i}$ lies in a collapsed stitch, in which case $u_{i}$ (resp. $v_{i}$) is its image under $\pi_{k}$ (resp. $\pi_{\infty}$). Let $\gamma_{k}$ be an $\varepsilon$–minimizing path in $X_{k}$ from $u_{1}$ to $u_{2}$ of length $\ell_{k}\leq d_{X_{k}}(u_{1},u_{2})+\varepsilon$. Lifting $\gamma_{k}$ to $\mathbb{E}^{N}$ produces a concatenation of Euclidean segments joining $a_{1}$ to $a_{2}$. Each time $\gamma_{k}$ crosses a stitch, the lift gains at most $\mathrm{diam}(S_{\alpha}^{(k)})\leq L$ in Euclidean length. Since distinct stitches are $\delta_{0}$–separated, any path of length $\leq 2R+2$ can meet at most $M=M(R,L,\delta_{0})$ of them, hence the lifted path has Euclidean length in $\ell_{k},\ \ell_{k}+ML]$. Performing the same construction in the limit space $X_{\infty}$ gives analogous bounds, and by (H1) the matched stitches are Hausdorff–close, so
$|d_{X_{k}}(u_{1},u_{2})-d_{X_{\infty}}(v_{1},v_{2})|\leq C(R,L,\delta_{0})\,\eta,$
where $\eta$ is the Hausdorff discrepancy of corresponding stitches. It follows that $\mathrm{dis}(\mathcal{R})\to 0$ as $k\to\infty$.
By (H1), each matched pair of stitches is Hausdorff–close by $\eta$ in $\mathbb{E}^{N}$. Combining this with the Euclidean triangle inequality and the previous length estimates gives an additive bound
$\big|d_{X_{k}}(u_{1},u_{2})-d_{X_{\infty}}(v_{1},v_{2})\big|\leq C(R,L,\delta_{0})\,\eta,$
where the constant $C(R,L,\delta_{0})$ depends only on the uniform local geometry and bounds the total contribution from the finitely many stitches that a minimizing path may cross. Choosing $\eta$ small enough so that $C\,\eta<\varepsilon$ yields $\mathrm{dis}(\mathcal{R})<\varepsilon$. Hence, for all sufficiently large $k$,
$d_{GH}\big(B_{R}^{X_{k}}(x_{k}),B_{R}^{X_{\infty}}(x_{\infty})\big)\leq\varepsilon.$
Finally, a standard diagonal argument over $R\to\infty$ gives
$(X_{k},x_{k})\xrightarrow[\;k\to\infty\;]{\mathrm{pGH}}(X_{\infty},x_{\infty}).$
[…]
3.2 Subsequential Gromov-Hausdorff convergence
The next question regarding Gromov-Hausdorff convergence is when a sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ has a Gromov-Hausdorff converging subsequence. Gromov’s pre-compactness theorem gives us sufficient conditions for a sequence of metric spaces to have such a subsequence, but the theorem requires a uniform upper bound on diameter and the number of disjoint balls of any given radius (intuitively, one should be able to fill the space with not too many small disjoint balls as one goes to smaller and smaller scales) [6]. Smocked spaces have unbounded diameter and do not satisfy these hypotheses, and so it is not clear if there should be a result on subsequential Gromov-Hausdorff convergence for these spaces.
In [2], it was shown that sequences of re-scalings of smocked spaces do converge in the pointed Gromov-Hausdorff sense, even though the spaces themselves do not satisfy the hypotheses of the Gromov pre-compactness theorem. There are some immediate examples of Gromov-Hausdorff converging subsequences which one can construct. For example, take a sequence of smocked spaces $\{X_{k}\}_{k=1}^{\infty}$ such that $X_{k}$ is a smocked space with one single smocking stitch for odd $k$ and such that $X_{k}$ is a smocked space with two smocking stitches which are the same compact subsets $\mathbb{E}^{n}$ for each even $k$. This sequence of smocked spaces contains a Gromov-Hausdorff converging subsequence given by $(X_{2k})$.
Generally speaking, we would like conditions for a sequence of smocked spaces to contain a pointed Gromov-Hausdorff converging subsequence. It turns out that to get such conditions it is sufficient to assume uniform global bounds on the smocking constants. Intuitively, the result says that one cannot pack in too many sets too closely together in the smocked space. This may explain why the examples found by Sormani et al. can be shown to have rescaled sequences which converge even though smocked spaces themselves do not meet the hypotheses of Gromov’s compactness theorem. However, to also have necessity, one must assume stronger local bounds on the smocking constants and local geometric control.
We begin by recalling Gromov’s precompactness theorem.
Proposition 3.4: Let $\{(Y_{\alpha},y_{\alpha})\}$ be a family of pointed proper metric spaces (i.e. all closed metric balls are compact). If for every $r,\varepsilon>0$ there is $N(r,\varepsilon)<\infty$ such that each $B_{r}^{Y_{\alpha}}(y_{\alpha})$ can be covered by at most $N(r,\varepsilon)$ balls of radius $\varepsilon$, then the family is precompact in the pointed Gromov–Hausdorff topology.
[…]
We next show that global bounds on the smocking constants are sufficient for Gromov-Hausdorff precompactness.
Proposition 3.5: Let $\{X_{k}\}$ be a sequence of smocked spaces obtained from $\mathbb{E}^{N}$ by collapsing collections of compact stitches, and let $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ be the smocking maps. Fix basepoints $x_{k}=\pi_{k}(0)$ and assume that there exists a set of smocking constants $L_{\max}<\infty$, $h<\infty$, $\delta>0$ and $L_{\min}>0$ independent of $k$ such that $L^{(k)}_{\text{max}}\leq L_{\text{max}}$, $h^{(k)}\leq h$, $\delta^{(k)}\geq\delta$, and $L^{(k)}_{\text{min}}\geq L_{\text{min}}$. Then the family $\{(X_{k},x_{k})\}$ is precompact in the pointed Gromov-Hausdorff topology (every sequence admits a subsequence which converges in the pointed Gromov-Hausdorff sense).
Proof: Fix $r>0$ and $\varepsilon>0$. The strategy of the proof is to produce a uniform bound $N(r,\varepsilon)$ on the number of disjoint $\varepsilon$-balls that can be packed into any ball $B_{r}^{X_{k}}(x_{k})$, with the required subsequence then following from Gromov’s precompactness theorem. By properties of the smocking map, we have for every $k$
$\pi_{k}^{-1}\!\big(B_{r}^{X_{k}}(x_{k})\big)\subset B_{r+L_{\max}}^{\mathbb{E}^{N}}(0).$
Since distinct stitches are separated by at least $\delta>0$ and every stitch has diameter $\leq L_{\max}$, only finitely many stitches intersect the compact Euclidean ball $B_{r+L_{\max}}^{\mathbb{E}^{N}}(0)$, with a bound which depends only on $r,L_{\max},\delta$ and $N$.
Next, set
$\varepsilon^{\prime}:=\min\{\varepsilon/4,\ \delta/4\}.$
Suppose that $\{B_{\varepsilon}^{X_{k}}(y_{i})\}_{i=1}^{M}$ is a collection of pairwise disjoint $\varepsilon$–balls contained in $B_{r}^{X_{k}}(x_{k})$. Choose lifts $v_{i}\in\pi_{k}^{-1}(y_{i})\subset B_{r+L_{\max}}^{\mathbb{E}^{N}}(0)$. We next claim that the Euclidean distance between any two distinct lifts satisfies $|v_{i}-v_{j}|\geq\varepsilon^{\prime}$. To see this, consider any two distinct points $y_{i},y_{j}\in X_{k}$ with $d_{X_{k}}(y_{i},y_{j})\geq\varepsilon$, and choose lifts $v_{i}\in\pi_{k}^{-1}(y_{i})$, $v_{j}\in\pi_{k}^{-1}(y_{j})$. There are two cases which must be checked separately.
In the first case, the lifts $v_{i},v_{j}$ lie in the same component of $\mathbb{E}^{N}\setminus\!\bigcup_{\alpha}S_{\alpha}^{(k)}$. $\pi_{k}$ is an isometry on that component, so we have $|v_{i}-v_{j}|_{\mathbb{E}^{N}}=d_{X_{k}}(y_{i},y_{j})\geq\varepsilon$. In the second case, the lifts lie in different components, hence they are separated by at least one collapsed stitch. Distinct stitches are $\delta$-separated in $\mathbb{E}^{N}$, so $|v_{i}-v_{j}|\geq\delta$. In particular, no two lifts of distinct $\varepsilon$–separated points in $X_{k}$ can be closer than $\varepsilon^{\prime}:=\min\{\varepsilon/4,\delta/4\}$ in $\mathbb{E}^{N}$. In both cases, we obtain $|v_{i}-v_{j}|\geq\varepsilon^{\prime}$ as claimed. It follows that the Euclidean balls $B_{\varepsilon^{\prime}/2}^{\mathbb{E}^{N}}(v_{i})$ are pairwise disjoint and lie inside the fixed compact set $B_{r+L_{\max}+\varepsilon^{\prime}/2}^{\mathbb{E}^{N}}(0)$.
By the Euclidean volume packing estimate,
$M\leq\frac{\operatorname{Vol}\big(B_{r+L_{\max}+\varepsilon^{\prime}/2}(0)\big)}{\operatorname{Vol}\big(B_{\varepsilon^{\prime}/2}(0)\big)}=:N(r,\varepsilon),$
which depends only on $r,\varepsilon,L_{\max},\delta$ and $N$ (in particular it is independent of $k$). As a result, the number of disjoint $\varepsilon$–balls that can be packed into $B_{r}^{X_{k}}(x_{k})$ is uniformly bounded by $N(r,\varepsilon)$.
[…]
Remark 3.6: Note that having uniform bounds on the smocking constants is a sufficient, but not a necessary, condition for subsequential convergence. For clarity, we will provide a counterexample in one dimension, but the argument can be easily generalized. For each integer $k\geq 1$, let $S_{k}=[k^{2},k^{2}+k]\subset\mathbb{R}$ and let $X_{k}$ be the smocked space obtained from $\mathbb{R}$ by collapsing $S_{k}$ to a single point. Denote by $\pi_{k}:\mathbb{R}\to X_{k}$ the smocking map and take the basepoint $x_{k}=\pi_{k}(0)$.
The maximal stitch diameter satisfies
$L_{\max}^{(k)}=\operatorname{diam}(S_{k})=k\longrightarrow\infty,$
so there is no global upper bound on $L_{\max}^{(k)}$. However, for any fixed radius $R>0$ the collapsed interval $S_{k}$ lies outside $\pi_{k}^{-1}\!\big(B_{R}^{X_{k}}(x_{k})\big)$ once $k$ is large, therefore the restriction of the metric on $B_{R}^{X_{k}}(x_{k})$ agrees with the Euclidean metric on [R,R][-R,R]. As a result, we have
$(X_{k},x_{k})\xrightarrow{\mathrm{pGH}}(\mathbb{R},0).$
This example shows that pointed Gromov-Hausdorff precompactness does not require uniform bounds on the smocking constants.
The assumption of global uniform bounds on the smocking constants is rather strong, since it controls the smocking pattern for all $k$ and across the entire space. In reality, pointed Gromov-Hausdorff convergence only sees geometry in neighborhoods around basepoints, so one may suspect that having bounded smocking constants should be both necessary and sufficient for pointed Gromov-Hausdorff convergence if the bounds are controlled in some local sense. The next proposition will focus on making this intuition precise.
Proposition 3.7: Let $\{(X_{k},x_{k})\}$ be a sequence of smocked spaces obtained from $\mathbb{E}^{N}$ by collapsing collections of compact stitches, with smocking maps $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ and basepoints $x_{k}=\pi_{k}(0)$. The sequence $\{(X_{k},x_{k})\}$ admits a subsequence which converges in the pointed Gromov–Hausdorff sense if and only if the following local uniform control holds: for every $r>0$, there exist smocking constants $L_{r}<\infty$, $\delta_{r}>0$ and an index $K_{r}$ such that for all $k\geq K_{r}$, every stitch that intersects $\overline{B_{r+L_{r}}^{\mathbb{E}^{N}}(0)}\;$ has diameter less than or equal to $L_{r}$ and any two distinct such stitches are separated by Euclidean distance at least $\delta_{r}$.
Proof: Begin by assuming that $\{(X_{k},x_{k})\}$ has a subsequence which converges in the pointed Gromov-Hausdorff sense. For simplicity, relabel and assume the whole sequence converges to a pointed limit $(X_{\infty},x_{\infty})$ (the argument below applies to any convergent subsequence). Fix $r>0$. By pointed Gromov-Hausdorff convergence, we have for every $\varepsilon>0$ a uniform packing bound $N(r,\varepsilon)$ for the balls $B_{r}^{X_{k}}(x_{k})$ for all sufficiently large $k$. Concretely, choose $\varepsilon<1$ and let $M:=N(r,\varepsilon)$. If for some large $k$ there were infinitely many distinct stitches intersecting $\overline{B_{r+1}^{\mathbb{E}^{N}}(0)}$, then by picking one point in each stitch and projecting to $X_{k}$ we would obtain arbitrarily many pairwise $\varepsilon$-separated points in $B_{r}^{X_{k}}(x_{k})$, contradicting the packing bound. It follows that for all large $k$ only finitely many stitches meet $\overline{B_{r+1}(0)}$, so the diameters of those stitches are uniformly bounded from above and their pairwise separations are uniformly bounded from below by positive constants. This gives the required $L_{r}$ and $\delta_{r}$ and an index $K_{r}$, so that the stated inequalities hold for all $k\geq K_{r}$.
For the other direction of the implication, begin by assuming that the local bounds exist for every $r>0$. Fix arbitrary $r>0$ and $\varepsilon>0$. By hypothesis there is $L_{r},\delta_{r}$ and an index $K_{r}$ such that for all $k\geq K_{r}$ every stitch meeting $\overline{B_{r+L_{r}}(0)}$ has diameter $\leq L_{r}$ and distinct such stitches are $\geq\delta_{r}$ apart. The argument in Proposition 3.3 can now be applied on the fixed radius $r$. Lift any collection of pairwise-disjoint $\varepsilon$-balls in $B_{r}^{X_{k}}(x_{k})$ to $\mathbb{E}^{N}$ and use the uniform separation and diameter bounds to produce a uniform Euclidean packing of radius $\varepsilon^{\prime}=\min\{\varepsilon/4,\delta_{r}/4\}$ in the fixed compact set $B_{r+L_{r}+\varepsilon^{\prime}/2}(0)$. It follows that there is a uniform bound $N(r,\varepsilon)$ on the number of disjoint $\varepsilon$-balls in $B_{r}^{X_{k}}(x_{k})$, valid for all $k\geq K_{r}$.
[…]
The previous proposition shows that that there is a geometric criterion for precompactness in terms of the smocking pattern: the constants $L_{r}$ and $\delta_{r}$ control, respectively, the maximal size of a stitch and the minimal Euclidean separation between distinct stitches inside a ball $B_{r+L_{r}}^{\mathbb{E}^{N}}(0)$. The smocking depth $h$ controls the geometry, but does not play a role in convergence and does not appear in the assumptions used for the proofs. The local bound hypothesis requires that smocking stitches should remain under uniform control within fixed bounded regions around the basepoint, although the geometry of the stitches may degenerate at arbitrarily large distances (stitches whose diameters tend to infinity whilst going to infinite separation do not affect the pointed limit). One may view the local bounds as a translation of the usual criteria needed for Gromov’s precompactness theorem into more easily visualizable data on the size and spacing of subsets of metric spaces.
3.3 Tangent cones at infinity
Finally, we show that for any finite dimensional normed vector space, there exists a smocked space whose unique tangent cone at infinity is that space. This answers in the affirmative an open question of Sormani et al. In general, there are examples of smocked spaces which do not have unique tangent cones at infinity, nor is the cone always a normed space [2].
Proposition 3.8: Let $(\mathbb{R}^{n},F)$ be a finite-dimensional normed vector space. Then there exists a smocked space $(X,d_{X})$ constructed in $\mathbb{E}^{n}$ and a basepoint $x_{0}\in X$ such that the rescaled pointed spaces $(X,\lambda^{-1}d_{X},x_{0})$ converge in the pointed Gromov-Hausdorff sense to $(\mathbb{R}^{n},F)$ as $\lambda\to\infty$. In particular, $(\mathbb{R}^{n},F)$ is the unique tangent cone at infinity of $(X,d_{X})$.
Proof: The tangent cone can be constructed explicitly using classical results on stable norms and homogenization [1, 7]. The first step is a standard one of approximating the ball by a rational polytope. Let $B_{F}=\{x:F(x)\leq 1\}$ be the unit ball of $F$. Since $B_{F}$ is convex and compact, there exists a symmetric finite subset $V=\{v_{1},\dots,v_{m}\}\subset\mathbb{Z}^{n}$ and weights $\ell_{i}:=F(v_{i})>0$ such that the polyhedral norm
$F_{V}(x):=\inf\Big\{\sum|a_{i}|\ell_{i}:\,x=\sum a_{i}v_{i}\Big\}$
satisfies $(1-\varepsilon)F(x)\leq F_{V}(x)\leq(1+\varepsilon)F(x)$ for all $x\in\mathbb{R}^{n}$ for arbitrarily small $\varepsilon>0$.
Let $L=\mathbb{Z}^{n}$ and define a graph $G=(V(G),E(G))$ with vertex set $V(G)=L$ and edges joining $p\in L$ to $p+v_{i}$ for every generator $v_{i}\in V$. Assign to the oriented edge $(p,p+v_{i})$ the length $\ell_{i}=F(v_{i})$. The induced shortest-path metric $d_{G}$ on $L$ is the weighted word metric associated to $V$. It is well known that the rescaled lattices $(L,\lambda^{-1}d_{G})$ converge in the pointed GH sense as $\lambda\to\infty$ to $(\mathbb{R}^{n},F_{V})$, and hence to $(\mathbb{R}^{n},F)$ because $F_{V}$ approximates $F$ uniformly.
Next, we must periodically embed this graph as a smocked subset of $\mathbb{E}^{n}$. This can be done as follows. For each vertex $p\in L$, insert a small closed ball $N_{p}:=\overline{B_{r}(p)}$ with radius $r>0$ chosen so that the balls (corresponding to nodes) are pairwise disjoint. For each edge $(p,p+v_{i})$, insert a narrow rectangular tube of Euclidean centerline length $\ell_{i}$ joining the boundaries $\partial N_{p}$ and $\partial N_{p+v_{i}}$. The tubes can be chosen to be sufficiently thin that distinct ones meet only at their endpoints. Let $S=\bigcup_{p\in L}N_{p}$ and define the smocked space $X:=\mathbb{E}^{n}/\!\!\sim$ by collapsing each node $N_{p}$ to a point via the smocking map $\pi:\mathbb{E}^{n}\to X$. This is a well-defined smocking of $\mathbb{E}^{n}$.
In the smocked space $X$, let $x_{p}=\pi(N_{p})$ for each vertex $p\in L$. For any two vertices $p,q\in L$, the distance between their images $x_{p}$ and $x_{q}$ equals the infimum of lengths of paths along connected tubes linking $N_{p}$ to $N_{q}$, since tubes are disjoint apart from their endpoints. This metric coincides with the weighted graph distance $d_{G}(p,q)$ up to an arbitrarily small error which can be eliminated by choosing the tube widths to be sufficiently small. It follows that the discrete subset $\{x_{p}\}_{p\in L}\subset X$ with induced metric $d_{X}$ is isometric to $(L,d_{G})$.
Since the construction is LL–periodic, the rescaled quotient metrics $(X,\lambda^{-1}d_{X})$ are periodic rescalings of the embedded weighted network. By the stable norm convergence and the fact that tube width can be neglected, it follows that
$(X,\lambda^{-1}d_{X},x_{0})\xrightarrow[\lambda\to\infty]{\mathrm{pGH}}(\mathbb{R}^{n},F),$
[…]
4 Smocked spaces with measures
4.1 Measured convergence of smocked measure spaces
We now extend the preceding convergence results to smocked spaces endowed with measures. A natural choice of measure on a smocked space $X=\mathbb{E}^{N}/\!\!\sim$ is the pushforward of Lebesgue measure under the smocking map $\pi:\mathbb{E}^{N}\to X$,
$\mu:=\pi_{\#}(\mathcal{L}^{N}),$
which gives the total Euclidean volume of preimages in $\mathbb{E}^{N}$. This defines a canonical metric measure structure $(X,d_{X},\mu)$ associated to every smocked space. We show below that the pointed Gromov–Hausdorff convergence results from Section 3 extend naturally to this measured setting.
Proposition 4.1: Let $(X_{k},x_{k},d_{k})$ be smocked spaces obtained from $\mathbb{E}^{N}$ by collapsing collections of compact stitches $\{S_{\alpha}^{(k)}\}$, and let $\pi_{k}:\mathbb{E}^{N}\to X_{k}$ be the associated smocking maps. Assume that $(X_{k},x_{k})\xrightarrow{\mathrm{pGH}}(X_{\infty},x_{\infty})$, where $X_{\infty}$ is obtained from the limiting stitches $\{S_{\beta}^{(\infty)}\}$. Define measures
$\mu_{k}:=(\pi_{k})_{\#}(\mathcal{L}^{N}),\qquad\mu_{\infty}:=(\pi_{\infty})_{\#}(\mathcal{L}^{N}).$
Then
$(X_{k},d_{k},\mu_{k},x_{k})\xrightarrow{\mathrm{pmGH}}(X_{\infty},d_{\infty},\mu_{\infty},x_{\infty})\quad\text{as }k\to\infty.$
Proof: By Proposition 3.5, for every $R>0$ there exist $\varepsilon_{k}\to 0$ and Borel $\varepsilon_{k}$–isometries $f_{k}:B_{R}^{X_{k}}(x_{k})\to B_{R+\varepsilon_{k}}^{X_{\infty}}(x_{\infty})$ with $f_{k}(x_{k})=x_{\infty}$ and $B_{R-\varepsilon_{k}}^{X_{\infty}}(x_{\infty})\subset(f_{k}(B_{R}^{X_{k}}(x_{k})))_{\varepsilon_{k}}$. Since the maps $\pi_{k}$ and $\pi_{\infty}$ differ from the identity on $\mathbb{E}^{N}$ only within unions of small stitches whose total Euclidean measure tends to zero, the pushforward measures $\mu_{k}$ converge weakly to $\mu_{\infty}$ on bounded sets. Equivalently, for every bounded continuous $\varphi$ on $B_{R}^{X_{\infty}}(x_{\infty})$,
$\int_{X_{k}}\!\varphi\!\circ f_{k}\,d\mu_{k}=\int_{\mathbb{E}^{N}}\!\varphi(f_{k}\!\circ\!\pi_{k})\,d\mathcal{L}^{N}\longrightarrow\int_{\mathbb{E}^{N}}\!\varphi(\pi_{\infty})\,d\mathcal{L}^{N}=\int_{X_{\infty}}\!\varphi\,d\mu_{\infty}.$
[…]
Corollary: If a sequence of smocking sets $S_{k}\subset\mathbb{E}^{N}$ converges in Hausdorff distance to $S_{\infty}$, and $(X_{k},d_{k})$, $(X_{\infty},d_{\infty})$ are the corresponding smocked spaces equipped with measures $\mu_{k}=(\pi_{k})_{\#}\mathcal{L}^{N}$ and $\mu_{\infty}=(\pi_{\infty})_{\#}\mathcal{L}^{N}$, then
$(X_{k},d_{k},\mu_{k},x_{k})\xrightarrow{\mathrm{pmGH}}(X_{\infty},d_{\infty},\mu_{\infty},x_{\infty}).$
This result shows that the smocking construction is stable under measured Gromov–Hausdorff convergence when equipped with the natural pushforward of Euclidean volume. In particular, the metric and measure structures both homogenize under the same geometric limits, confirming that smocked spaces extend consistently to the metric measure setting.
5 Conclusions
Smocked spaces, introduced by Sormani et al. as a geometric model for “stitched” or “collapsed” Euclidean structures, form a class of quotient spaces whose analytic and geometric properties have mostly been unexplored. In this work, we have provided a systematic treatment of their convergence theory and answered several open questions regarding their convergence properties. In particular, we have:
  • established the conditions under which Hausdorff convergence of the smocking sets implies pointed Gromov-Hausdorff convergence of the associated smocked spaces
  • identified local uniform bounds on the smocking constants as sufficient and necessary for Gromov–Hausdorff precompactness
  • constructed examples showing that global or uniform bounds are not required, thereby clarifying the geometric mechanisms underlying previous examples of convergence
  • proved that every finite-dimensional normed vector space arises as a tangent cone at infinity of a smocked space
  • extended the convergence theory to the setting of smocked metric measure spaces, showing stability under pointed measured Gromov–Hausdorff convergence.
These results shows that smocked spaces have a rigorous convergence theory similar to other established quotient space constructions in metric geometry, such as Alexandrov spaces and length-space gluings. The ’smocking’ concept provides a flexible mechanism for creating non-trivial limits of Euclidean spaces via controlled local collapse and hence these spaces may be a useful model for testing questions on metric convergence and tangent cones in a very geometrically explicit setting. They may also be useful as simplified models for metric collapse, where analytic and geometric quantities interact but can still be computed using elementary means.
The results developed here suggest several directions for future research. These include studying curvature or rectifiability properties of smocked spaces, formulating analogues of RCD-type or synthetic curvature conditions, and developing a corresponding theory of Gromov–Wasserstein convergence. In summary, this work demonstrates that smocked spaces are a non-trivial construction. Under appropriate geometric control, they exhibit rich convergence behavior and capture a wide variety of limit geometries. It would be interesting to investigate whether the smocking construction extends naturally to Wasserstein or Gromov–Wasserstein spaces, and whether analogues of contractivity or curvature dimension conditions hold for such smocked Wasserstein metrics.
Acknowledgments
This research was partly conducted whilst the author was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP).
Funding
This study was not funded.
Conflict of Interest
The author declares that he has no conflict of interest.
Ethical approval
This article does not contain any studies with human participants or animals performed by any of the authors.